OpenStudy (anonymous):

HELP! What would be the first step in solving this....6/7x+2=18

OpenStudy (♪chibiterasu):

Well first we just subtract 2 from both sides: 6/7x + 2 - 2 = 18 - 2 6/7x = 16 What do you think we do next?

OpenStudy (♪chibiterasu):

We subtract 2 from both sides since it is added on one side. We subtract to cancel it out.

OpenStudy (♪chibiterasu):

And now I'm alone >_>